# Advancing Model-Assisted Design in Phase I Trials through Bayesian Dose-Response Model

### v1

@@ 5 Discussion for reference dose and prior distribution
Tables 5, 6, and 7 show the results obtained using logit, log-log, and clog-log link functions, respectively. In scenario 2, the accuracy of MTD selection declines as the reference dose increased, while in Scenario 3, it shows improvement. For the remaining scenarios, the reference dose has little impact. A common feature of scenarios 2 and 3, which exhibit significant variability in results, is that the MTDs are set at higher dose levels. Table 8 shows the averages of $n_{j}$ and $m_{j}$ over 1000 datasets for each dose level in every scenario. Note that $n_{j}$ and $m_{j}$ remain unchanged regardless of the reference dose. In scenarios 2 and 3, the number of DLTs, $m_{j}$, is lower than in other scenarios. This is because, in trials where the MTD is set at a higher dose, there is typically a lower number of DLTs at higher doses. Generally, in scenarios where an insufficient number of DLTs are observed, such as when the MTD is near the highest dose, the influence of the likelihood on DLT probability estimation decreases, making the prior distribution relatively more influential. Consequently, the dependence of the prior distribution on the reference dose setting may have led to increased variability in the results of scenarios 2 and 3, particularly due to a few number of DLTs at higher doses. In contrast, in scenario 4, where the number of DLT occurrences is larger, the influence of the reference dose selection is minimal. Nevertheless, since cases where the MTD is set at the highest dose are extremely rare in Phase I oncology trials, the observed instability in scenarios 2 and 3 is not of major concern.
Table 5: The proportion of MTD selections across 1,000 simulations in each scenario for the PAVA-based method and the proposed method using the logit model when the reference dose $d^{*}$ is $d_{4}$, $d_{5}$, or $d_{6}$. The optimal MTD is highlighted in bold.
@@
Table 6: The proportion of MTD selections across 1,000 simulations in each scenario for the PAVA-based method and the proposed method using the log-log model when the reference dose $d^{*}$ is $d_{4}$, $d_{5}$, or $d_{6}$. The optimal MTD is highlighted in bold.
@@
Table 7: The proportion of MTD selections across 1,000 simulations in each scenario for the PAVA-based method and the proposed method using the clog-log model when the reference dose $d^{*}$ is $d_{4}$, $d_{5}$, or $d_{6}$. The optimal MTD is highlighted in bold.

### v2

@@ 5 Discussion for reference dose and prior distribution
Tables 6, 7, and 8 show the results obtained using logit, log-log, and clog-log link functions, respectively. In scenario 2, the accuracy of MTD selection declines as the reference dose increased, while in Scenario 3, it shows improvement. For the remaining scenarios, the reference dose has little impact. A common feature of scenarios 2 and 3, which exhibit significant variability in results, is that the MTDs are set at higher dose levels. Table 9 shows the averages of $n_{j}$ and $m_{j}$ over 1000 datasets for each dose level in every scenario. Note that $n_{j}$ and $m_{j}$ remain unchanged regardless of the reference dose. In scenarios 2 and 3, the number of DLTs, $m_{j}$, is lower than in other scenarios. This is because, in trials where the MTD is set at a higher dose, there is typically a lower number of DLTs at higher doses. Generally, in scenarios where an insufficient number of DLTs are observed, such as when the MTD is near the highest dose, the influence of the likelihood on DLT probability estimation decreases, making the prior distribution relatively more influential. Consequently, the dependence of the prior distribution on the reference dose setting may have led to increased variability in the results of scenarios 2 and 3, particularly due to a few number of DLTs at higher doses. In contrast, in scenario 4, where the number of DLT occurrences is larger, the influence of the reference dose selection is minimal. Nevertheless, since cases where the MTD is set at the highest dose are extremely rare in Phase I oncology trials, the observed instability in scenarios 2 and 3 is not of major concern.
Given that the influence of reference dose selection is most pronounced in scenarios where the MTD is at higher dose levels (which are extremely rare in Phase I oncology trials), we recommend a conservative approach of selecting the reference dose at or near the middle of the dose range (e.g., $d_{3}$ or $d_{4}$ in a 6-dose study). This minimizes potential instability while maintaining good performance across most realistic scenarios. For investigators concerned about reference dose sensitivity, conducting a sensitivity analysis with different reference dose choices during the trial planning phase would provide additional assurance about the robustness of the design.
Table 6: The proportion of MTD selections across 1,000 simulations in each scenario for the proposed method using the logit model when the reference dose $d^{*}$ is $d_{4}$, $d_{5}$, or $d_{6}$. The optimal MTD is highlighted in bold.
@@
Table 7: The proportion of MTD selections across 1,000 simulations in each scenario for the proposed method using the log-log model when the reference dose $d^{*}$ is $d_{4}$, $d_{5}$, or $d_{6}$. The optimal MTD is highlighted in bold.
@@
Table 8: The proportion of MTD selections across 1,000 simulations in each scenario for the proposed method using the clog-log model when the reference dose $d^{*}$ is $d_{4}$, $d_{5}$, or $d_{6}$. The optimal MTD is highlighted in bold.